# PFGuard: A Generative Framework with Privacy and Fairness Safeguards

### v1

@@ Appendix D Experimental Settings
Continuing from Sec. 5, we provide more details on experiment settings. In all experiments, we use PyTorch and perform experiments using NVIDIA Quadro RTX 8000 GPUs. Also, we repeat all experiments 10 times and report the mean and standard deviation of the top 3 results. The reason we report the top-3 results is to favor the simple privacy-fairness baselines (e.g., “Reweighting” in Table 5.2), which tend to fail frequently. We compare their best performances with PFGuard.
@@ D.1 Datasets and Bias Settings
Continuing from Sec. 5, we provide more details on datasets. We use three datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CelebA (Liu et al., 2015). MNIST and FashionMNIST contain grayscale images with 28 x 28 pixels and 10 classes. Both datasets have 60,000 training examples and 10,000 testing examples. CelebA contains 202,599 celebrity face images. We use the official preprocessed version with face alignment and follow the official training and testing partition (Liu et al., 2015). Note that we are using image datasets instead of the traditional smaller tabular benchmarks for fairness because our goal is to make PFGuard work on higher dimensional data such as images.
@@
  • Binary Class Bias. For MNIST, we set digit “3” as the majority class $Y=1$ and “1” as the minority class $Y=0$; for FasionMNIST, we set “Sneakers” as $Y=1$ and “Trousers” as $Y=0$. For each class pair, we select two classes that share the fewest false negatives and thus can be considered independent, following the convention of prior approaches (Bagdasaryan et al., 2019; Farrand et al., 2020; Ganev et al., 2022). We set bias level as 2, meaning the minority class $Y=0$ is 2 times smaller than the majority class $Y=1$. After creating bias, we apply random affine transformations to augment the datasets, ensuring they match the original dataset size.
  • Multi-class Bias. We set “8” as the minority class $Y=0$, reducing its size while maintaining the size of other 9 classes, following the above prior approaches. We vary the bias level from 1 to 10.
  • Subgroup Bias. For both MNIST and FashionMNIST datasets, we use image rotation to define subgroups. We set non-rotated images as the majority group $S=1$ and rotated image as the minority group $S=0$. We also considered other options including adding lines and changing colors, but we observed that the other options often show the adverse affect of making the images noisier and thus reducing the model accuracy unnecessarily. The rotation also allows for simple and effective verification of subgroup labels in generated synthetic data by comparing the mean values of synthetic image vectors to the centroids of real image vectors. To validate this heuristic, we compared the results with 400 manually labeled images from each baseline model and observed high accuracy (e.g., 96.5% for MNIST).
  • Unknown Subgroup Bias. In the previous subgroup bias setting, $S$ labels are not used during model training; they are only used for evaluation purposes after training.
@@
We create binary class bias using gender attributes, where we set female and male images as $Y=1$ and $Y=0$, respectively. As discussed in the main text, DP generative models often show low performance on CelebA in challenging bias scenarios like multi-class bias, which can hinder the reliability of fairness analyses (e.g., a random generator achieves perfect fairness by outputting random images regardless of data groups). Notably, we show that DP generative models can produce highly biased synthetic data even in this simple binary class bias setting (Table 4).
@@ D.2 Baselines
We use three state-of-the-art PTEL-based generative models: GS-WGAN (Chen et al., 2020), G-PATE (Long et al., 2021), and DataLens (Wang et al., 2021a). For all models, we use their official Github codes to implement their models and to use their best-performing hyperparameters for MNIST, FashionMNIST and CelebA.
  • GS-WGAN. GS-WGAN is extensively used in our experiments, as it leverages both PTEL and DP-SGD (Abadi et al., 2016) to ensure DP and thus allows various integration with other techniques. Their approach first trains a multiple teacher ensemble and considers one teacher as the representative of the other teachers. The output of representative teacher (i.e., gradients) is then sanitized with a DP-SGD based mechanism to train a DP generator. Compared to DP-SGD which operates on the whole minibatch, the DP mechanism in GS-WGAN operates on each data sample and thus can be considered as a composition of $B$ Gaussian mechanism where $B$ is the minibatch size. Our fair sampling preserves their sensitivity analyses despite the potential oversampling as it does not change the sensitivity of each Gaussian mechanism on one input data (i.e, $2C$ due to triangle inequality).
  • G-PATE and DataLens. G-PATE and DataLens leverage teachers’ votes on intermediate gradients to update the generator. To generate histograms of teachers’ votes and sanitize with DP mechanisms, G-PATE uses random projection and gradient discretization while DataLens uses a top-k stochastic sign quantization of the gradients. Our fair sampling preserves their sensitivity analyses despite the potential oversampling as each teacher still throws only one vote.
@@
We use a prominent fair training approach based on reweighting (Choi et al., 2020), and use two recent classification techniques which address both privacy and fairness: DP-SGD-F (Xu et al., 2020) and DPSGD-Global-Adapt (Esipova et al., 2022).
  • Reweighting. As outlined in Sec. 3, the reweighting approach modifies the loss term of a generative model to achieve fairness. We use likelihood ratio computed in each bias setting as the reweighting factor and only modifies the loss term of a discriminator, following their approach. When computing likelihood ratio, we directly compute the value as in Eq. 5 using sensitive group labels for subgroup bias setting; we estimate the value using binary classification approach, implemented in their official Github code for unknown subgroup setting. We note that a public reference dataset is required in the estimation process to effectively train a binary classifier.
  • DP-SGD-F and DPSGD-Adapt-Global DP-SGD-F and DPSGD-Adapt-Global are both fair variants of DP-SGD, where clipping bounds are dynamically adjusted to control the fairness-utility tradeoff. To prevent excessive gradient clipping for minority data group samples, DP-SGD-F employs a groupwise clipping approach where each data group has its own clipping bound, while DPSGD-Adapt employs a scaling approach where all per-sample gradients are scaled down depending on a dynamically adjusted scaling factor. As DP-SGD-F and DPSGD-Adapt-Global do not provide the official codes to our knowledge, we faithfully implemented each algorithm based on their papers. We note that DP-SGD-F is not applicable in the unknown subgroup setting as they require number of group samples present in the batch to compute clipping bounds for each group; DPSGD-Global-Adapt is applicable as the scaling factor does not require knowledge on group labels.
@@ Appendix E Additional Experiments
Continuing from Sec. 5, Sec. 5.1, Sec. 5.3, and Sec. 5.4, we provide more experimental results.
E.1 Comparison of Computational Time
Continuing from Sec. 5.3, we compare the computational time when integrating PFGuard with existing PTEL-based generative models. Table shows that PFGuard incurs minimal overhead in computational time ($<4\%$), due to the simple modification in minibatch sampling for fairness.
@@ E.1 Comparison of Computational Time
E.2 Experimental Results on FashionMNIST
@@
Continuing from Sec. 5.3, we show the results of the analysis in synthetic data (Table 6) and downstream tasks (Table 7 ) evaluated on FashionMNIST. Compared to the results evaluated on MNIST, private generative models often generate more imbalanced synthetic data w.r.t. sensitive groups and exhibit lower overall image quality. In comparison, PFGuard consistently improves both fairness and overall utility in most cases, similar to the results observed in the MNIST evaluation.
@@
E.3 Additional Normalization Technique for Faster Convergence
Continuing from Sec. 5.3, we investigate the impact of the normalization factor on the overall image quality of PFGuard. While we use a traditional normalization factor $N_{1}=\sum_{i}h(x_{i})$ for $w(x_{i})\propto h(x_{i})$ in our SIR-based sampling algorithm, we can employ additional normalization techniques to boost the performance. For example, we can use $N_{2}=\sum_{i}h(x_{i})/N_{-i}$ for $w(x_{i})\propto h(x_{i})/N_{-i}$, where $N_{-i}=\sum_{i}h(x_{i})-h(x_{i})$, which is known to help faster convergence of SIR algorithms to the target distribution $p_{\text{bal}}$ (Skare et al., 2003).
We thus compare the overall image quality resulting from two different normalization options, $N_{1}$ and $N_{2}$, varying batch sizes to analyze their effects on model convergence. We note that using a larger batch size can change the DP analysis (i.e., the more a model uses the data, the weaker privacy it provides). To effectively compare the difference in convergence speed, we create both binary class bias and subgroup bias on the MNIST dataset, where class $Y=0$ is 3 times smaller than class $Y=1$, and group $S=0$ is 3 times smaller than group $S=1$.
Table 8 shows the comparison of image quality using the MNIST dataset, measuring image quality with FID, where a lower value is better. While both $N_{1}$ and $N_{2}$ demonstrate comparable performance when using a large batch size, the performance gap becomes more evident as the batch size decreases. This empirical evidence shows that the performance of PFGuard can be further improved by additionally employing various normalization techniques.
Table 8: Influence of normalization factor on MNIST with corresponding DP guarantees ($\varepsilon$) to different batch sizes. GS-WGAN is used as the base DP generator.
Normalization factor
Batch size $N_{1}$ $N_{2}$
128 ($\varepsilon=29.91$) 75.05±2.26 74.58±2.96
64 ($\varepsilon=19.58$) 75.35±6.67 72.20±4.58
32 ($\varepsilon=9.99$) 82.68±7.13 78.18±1.85
E.4 Full Results with Standard Deviation
Continuing from Sec. 5.2 and Sec. 5.4, we show full results with standard deviation. Table 9 and Table 10 shows the full results of Table 4.

### v2

@@ Appendix D Experimental Settings
Continuing from Sec. 5, we provide more details on experimental settings. In all experiments, we use PyTorch and perform experiments using NVIDIA Quadro RTX 8000 GPUs. Also, we repeat all experiments 10 times and report the mean and standard deviation of the top 3 results. The reason we report the top-3 results is to favor the simple privacy-fairness baselines (e.g., “Reweighting” in Table 5.2), which we observe to fail frequently. We compare their best performances with PFGuard.
@@ D.1 Datasets and Bias Settings
Continuing from Sec. 5, we provide more details on datasets. We use three datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CelebA (Liu et al., 2015). MNIST and FashionMNIST contain grayscale images with 28 x 28 pixels and 10 classes. Both datasets have 60,000 training examples and 10,000 testing examples. CelebA contains 202,599 celebrity face images. We use the official preprocessed version with face alignment and follow the official training and testing partition (Liu et al., 2015). We mainly use image datasets instead of the traditional smaller tabular benchmarks for fairness because our goal is to make PFGuard work on higher dimensional data such as images. Additional results on tabular datasets are provided in Sec. E.2.
@@
  • Binary Class Bias. For MNIST, we set digit “3” as the majority class ${\mathbf{y}}=1$ and “1” as the minority class ${\mathbf{y}}=0$. For FasionMNIST, we set “Sneakers” as ${\mathbf{y}}=1$ and “Trousers” as ${\mathbf{y}}=0$. For each class pair, we select two classes that share the fewest false negatives and thus can be considered independent, following the convention of prior approaches (Bagdasaryan et al., 2019; Farrand et al., 2020; Ganev et al., 2022). We set bias level as 2, which means the minority class ${\mathbf{y}}=0$ is 2 times smaller than the majority class ${\mathbf{y}}=1$. After creating bias, we apply random affine transformations to augment the datasets to match the original dataset size.
  • Multi-class Bias. We set “8” as the minority class ${\mathbf{y}}{=}0$ and reduce its size while maintaining the size of the other 9 classes, following the aforementioned prior approaches. We vary the bias level from 1 to 10.
  • Subgroup Bias. For both MNIST and FashionMNIST datasets, we use image rotation to define subgroups. We set non-rotated images as the majority group ${\mathbf{s}}=1$ and rotated image as the minority group ${\mathbf{s}}=0$. We also considered other options including adding lines and changing colors, but we observed that the other options often show the adverse affect of making the images noisier and thus reducing the model accuracy unnecessarily. The rotation also allows for simple and effective verification of subgroup labels in generated synthetic data by comparing the mean values of synthetic image vectors to the centroids of real image vectors. To validate this heuristic, we compared the results with 400 manually labeled images from each baseline model and observed high accuracy (e.g., 96.5% for MNIST).
  • Unknown Subgroup Bias. In the previous subgroup bias setting, ${\mathbf{s}}$ labels are not used during model training and only used for evaluation purposes after training.
@@
We create binary class bias using gender attributes, where we set female and male images as ${\mathbf{y}}=1$ and ${\mathbf{y}}=0$, respectively. As discussed in the main text, DP generative models often show low performance on CelebA in challenging bias scenarios like multi-class bias, which can hinder the reliability of fairness analyses (e.g., a random generator achieves perfect fairness by outputting random images regardless of data groups). Notably, we show that DP generative models can produce highly biased synthetic data even in this simple binary class bias setting (Table 4).
@@ D.2 Baselines
We use three state-of-the-art PTEL-based generative models: GS-WGAN (Chen et al., 2020), G-PATE (Long et al., 2021), and DataLens (Wang et al., 2021a). For all models, we refer to their official Github codes to implement their models and to use their best-performing hyperparameters for MNIST, FashionMNIST, and CelebA.
  • GS-WGAN. GS-WGAN is widely used in our experiments, as it leverages both PTEL and DP-SGD (Abadi et al., 2016) to ensure DP and thus allows for various integration with other techniques. Their approach first trains a multiple teacher ensemble and considers one teacher as the representative of the other teachers. Output of the representative teacher (i.e., gradients) is then sanitized with a DP-SGD based mechanism to train a DP generator. Compared to DP-SGD which operates on the whole minibatch, their DP mechanism operates on each data sample and thus can be considered as a composition of $B$ Gaussian mechanism where $B$ is the minibatch size. Our fair sampling preserves their sensitivity analyses despite the potential oversampling, as it does not change the sensitivity of each Gaussian mechanism on one input data (i.e., $2C$ due to the triangle inequality).
  • G-PATE and DataLens. G-PATE and DataLens leverage teachers’ votes on intermediate gradients to update the generator. To sanitize these teachers’ votes with DP mechanisms, G-PATE uses a random projection and a gradient discretization, while DataLens uses a top-k stochastic sign quantization. Our fair sampling preserves their sensitivity analyses despite the potential oversampling, as each teacher still can throw only one vote.
@@
We use 1) a prominent fair training approach based on reweighting (Choi et al., 2020) and 2) two recent classification techniques which address both privacy and fairness: DP-SGD-F (Xu et al., 2020) and DPSGD-Global-Adapt (Esipova et al., 2022).
  • Reweighting. As outlined in Sec. 3, the reweighting approach computes the likelihood ratio and modifies the loss term of a discriminator to achieve fairness. For subgroup bias setting, we directly compute the likelihood ratio using sensitive group labels (Eq. 5). For unknown subgroup setting, we estimate the value using binary classification approach with their official Github code. During this estimation process, a public reference dataset is required to effectively train a binary classifier.
  • DP-SGD-F and DPSGD-Adapt-Global. DP-SGD-F and DPSGD-Adapt-Global are fair variants of DP-SGD where clipping bounds are dynamically adjusted to control the fairness-utility tradeoff. To prevent excessive gradient clipping for minority data group samples, DP-SGD-F employs a groupwise clipping approach where each data group has its own clipping bound, while DPSGD-Adapt employs a scaling approach where all per-sample gradients are scaled down depending on a dynamically adjusted scaling factor. As DP-SGD-F and DPSGD-Adapt-Global do not provide the official codes to our knowledge, we faithfully implemented each algorithm based on their papers. We note that DP-SGD-F is not applicable in the unknown subgroup setting, as they require number of subgroup samples present in the batch to compute clipping bounds for each subgroup; in contrast, DPSGD-Global-Adapt is applicable in the unknown subgroup setting, as the scaling factor does not require knowledge on subgroup labels.
@@ Appendix E Additional Experiments
Continuing from Sec. 5,  5.15.3, and  5.4, we provide more experimental results.
E.1 More experiments on Privacy-Fairness-Utility Tradeoff
Continuing from Sec. 5.2, we provide more results on the privacy-fairness-utility tradeoff.
Fairness-Utility Tradeoff with Varying Privacy Levels
We analyze the fairness-utility tradeoff under varying privacy levels, following (Tran et al., 2021b). Specifically, we examine how training with PFGuard impacts the original fairness-utility tradeoff of underlying private generative models, using the MNIST dataset under the subgroup bias setting. The results are shown in Fig. 8 and Fig. 8. In weaker privacy regimes (i.e., higher values of $\varepsilon$), both the privacy-only generative model and PFGuard converge to similar performance levels. However, PFGuard show notable fairness improvements in stronger privacy regimes, albeit it also shows utility degradation due to slower convergence in the early training stages. We suspect this slower convergence is a natural consequence of learning a more balanced distribution, since the generative model should capture more diverse input features at the early training stage and thus can be more challenging.
Figure 6: Visualization of Pareto frontier results of PFGuard. A darker color indicates a more stronger privacy constraint. Both fairness metric (Distributon disparity) and utility metric (Overall FID) are lower the better.
Figure 7: Fairness-utility tradeoff of GS-WGAN (i.e., private generative model).
Figure 8: Fairness-utility tradeoff of GS-WGAN when trained with PFGuard.
Pareto Frontier Results
Fig. 8 visualizes the Pareto frontier results of PFGuard, showing the privacy-fairness-utility tradeoff. The nonlinear surface highlights the intricate relationship between these three objectives. As privacy constraints weaken (i.e., the light-colored region), both utility and fairness improve, converging toward a more favorable region. However, as privacy constraints become stronger, the utility narrows down to a specific range, while fairness greatly varies, which is particularly evident in the purple-colored region. This suggests that stronger DP noise consistently degrades the image quality, but results in highly variable fairness outcomes, which implies that there can be sweet-spot regions that can achieve both fairness and utility.
E.2 Experimental results on Tabluar Data
Continuing from Sec. 5.3, we provide additional results with a tabular dataset. Although tabular data are not of our immediate interest given our focus on scalability to high-dimensional data such as images, we show how PFGuard can support tabular data as well as image data.
Experimental Setups
[Dataset] We use Adult (Kohavi, 1996), which contains 43,131 examples of demographic information of individuals along with a binary label of whether their annual income is greater than 50k. We use gender as the sensitive attribute ${\mathbf{s}}$ and income as the class label ${\mathbf{y}}$. [Baseline] We use DP-WGAN (Xie et al., 2018) and PATE-GAN (Jordon et al., 2018) for private-only generative models, and FFPDG (Xu et al., 2021) for both fair and private generative models. Since PATE-GAN is a PTEL-based generative model, we use PATE-GAN as the base generator of PFGuard. We note that FFPDG is the most closest work of ours, which mainly focuses on tabular data and poses challenges when scaling to images – see more details in Sec. 6. [Metric] We use privacy and fairness metrics from the main text and AUROC as the utility metric, following FFPDG for effective comparison.
Privacy-Fairness-Utility Tradeoff
Table E.2 shows the privacy-fairness-utility performances on the Adult dataset. Notably, PFGuard achieves fairness comparable to fair-only generative models while maintaining comparable utility with privacy-only generative models and FFPDG. In addition, Fig. 10 shows the Pareto frontier results compared to privacy-only generative models, where PFGuard outperforms other baselines by showing curves closer to the ideal region (i.e., low fairness discrepancy and high utility). These results on tabular data are consistent with the results on image data, highlighting the effectiveness and flexibility of PFGuard to a wide range of real-world applications by supporting both tabular and image datasets.
Table 5: Comparison of privacy-fairness-utility performance on Adult under $\varepsilon{=}1$, using PATE-GAN as the base DP generator. The first five rows represent upper bound performances for vanilla, DP-only, and fair-only models. Evaluations cover subgroup bias. Lower values are better across all metrics except AUROC. We boldface the best results and underline the second best results.
Privacy Fairness Utility
Method $\varepsilon~{}~{}~{}$ EO Disp. ($\downarrow$) Dem. Disp. ($\downarrow$) AUROC ($\uparrow$)
Vanila 0.56 0.58 0.80
Fair-only 0.07 0.07 0.75
DP-WGAN 0.31 0.30 0.69
PATE-GAN 0.19 0.22 0.74
RON-Gauss 0.18 0.14 0.70
FFPDG 0.12 0.20 0.75
[…]
@@ E.2 Experimental results on Tabluar Data
Figure 9: Pareto frontier results of DP generative models evaluated on tabular dataset (Adult). Upper left region denotes ideal case with low fairness discrepancy (EO Disparity) and high utility (AUROC).
Figure 10: FID values with different normalization factors, evaluated on MNIST using different batch sizes. GS-WGAN is used as the base DP generator. $N_{1}$ denotes a default normalization factor (Sec. E.4); $N_{2}$ denotes normalization factor from (Skare et al., 2003).
Normalization factor
Batch size $N_{1}$ $N_{2}$
128 ($\varepsilon=29.91$) 75.05±2.26 74.58±2.96
64 ($\varepsilon=19.58$) 75.35±6.67 72.20±4.58
32 ($\varepsilon=9.99$) 82.68±7.13 78.18±1.85
E.3 Comparison of Computational Time
Continuing from Sec. 5.3, we compare the computational time when integrating PFGuard with existing PTEL-based generative models. Table 6 shows that PFGuard incurs minimal overhead in computational time ($<4\%$), due to the simple modification in minibatch sampling for fairness.
E.4 Additional Normalization Technique for Faster Convergence
Continuing from Sec. 5.3, we investigate the impact of the normalization factor on the overall image quality of PFGuard. While we use a traditional normalization factor $N_{1}=\sum_{i}h(x_{i})$ for $w(x_{i})\propto h(x_{i})$ in our SIR-based sampling algorithm, we can employ additional normalization techniques to boost the performance. For example, we can use $N_{2}=\sum_{i}h(x_{i})/N_{-i}$ for $w(x_{i})\propto h(x_{i})/N_{-i}$ where $N_{-i}=\sum_{i}h(x_{i})-h(x_{i})$, which is known to help faster convergence of SIR algorithms to the target distribution $p_{\text{bal}}$ (Skare et al., 2003). To investigate the influence on model convergence of normalization factors, we compare the overall image quality resulting from two different normalization factors $N_{1}$ and $N_{2}$ with varying batch sizes. We create both binary class bias and subgroup bias in the MNIST dataset with bias level 3 (see Sec. D.1 for details) to make a more challenging setting to effectively compare convergence speeds. We note that using a larger batch size can change the corresponding DP guarantee, as using large batches leads to more data usage.
Table 10 shows the comparison of image quality measured with FID, where a lower value indicates better quality. Although both $N_{1}$ and $N_{2}$ demonstrate comparable performance when using a large batch size, the performance gap becomes more evident as the batch size decreases. This empirical evidence shows that the performance of PFGuard can be further improved by additionally employing various normalization techniques.
E.5 Comparison with a Different Sampling Strategy
Continuing from Sec. 5.2, we provide a performance comparison with a different sampling scenario to enhance fairness. If sensitive attribute labels are available, one natural strategy is to train teachers exclusively on each data group to learn group representations, which is a commonly used strategy in stratified sampling approaches (Meng, 2013). We compare this stratified sampling strategy against PFGuard’s importance sampling strategy, which trains teachers on randomly partitioned data, assessing their respective privacy-fairness-utility tradeoffs.
Table 7 shows that while stratified sampling achieves comparable fairness to importance sampling, it significantly reduces utility. A potential theoretical explanation is the downgraded utility of the teacher ensemble. Training teachers on non-i.i.d. datasets can lead to inconsistent convergence across teachers, leading to low consensus during teacher voting (Dodwadmath & Stich, 2022). This low consensus not only reduces the ensemble’s prediction accuracy (i.e., utility), but also increases privacy costs during DP aggregation (Papernot et al., 2018) – ultimately leading to worse privacy-fairness-utility tradeoff. These results suggest that while training teachers on heterogeneous data may be useful in other contexts, importance sampling is better suited for ensemble learning context to achieve both privacy and fairness.
Table 7: Comparison of privacy-fairness-utility performance of sampling strategies, evaluated on MNIST under the subgroup bias setting. GS-WGAN is used as the base DP generator. Lower values are better across all metrics. We boldface the best results and underline the second best results.
Privacy Fairness Utility
Method $\varepsilon~{}~{}~{}$ KL ($\downarrow$) Dist. Disp. ($\downarrow$) FID ($\downarrow$)
DP-only 10 0.177 0.383 77.97
Stratified sampling 10 0.090 0.209 135.35
Importane sampling 10 0.067 0.242 83.67
E.6 Extension with other fairness techniques
Continuing from Sec. 5.3 and Sec. C.3, we show experimental results on supporting other fairness techniques. In particular, we use GOLD (Mo et al., 2019), which aims to achieve Rawlsian Max-Min fairness (Joseph et al., 2016) and is compatible with PFGuard (see Sec. C.3 for more details).
Table 8 shows the privacy-fairness-utility performance when applying GOLD in the unknown sensitive attribute setting. GOLD achieves the best performance for the smallest (i.e., worst-case) group, aligning with its goal of Rawlsian Max-min fairness. However, GOLD does not surpass PFGuard in group fairness metrics or overall utility. These results suggest that incorporating additional fairness techniques can provide flexibility for different use cases, where GOLD can be particularly beneficial in applications prioritizing Rawlsian Max-min fairness over group fairness.
Table 8: Comparison of privacy-fairness-utility performance of additional fairness technique (GOLD (Mo et al., 2019)), evaluated on MNIST under unknown subgroup bias setting. GS-WGAN is used as the base DP generator. Lower values are better across all metrics. We boldface the best results and underline the second best results.
Privacy Fairness Utility
Method $\varepsilon~{}~{}~{}$ KL ($\downarrow$) Dist. Disp. ($\downarrow$) Smallest group FID ($\downarrow$) FID ($\downarrow$)
DP-only 10 0.177 0.383 101.39 77.97
PFGuard 10 0.004 0.041 89.43 89.76
PFGuard + GOLD 10 0.090 0.209 84.52 100.39
E.7 Experimental Results on FashionMNIST
@@
Continuing from Sec. 5.3, we show PFGuard’s performances in synthetic data (Table 9) and downstream tasks (Table 10) evaluated on the FashionMNIST dataset. Compared to the results evaluated on MNIST, private generative models often generate more imbalanced synthetic data w.r.t. sensitive groups and exhibit lower overall image quality. In comparison, PFGuard consistently improves both fairness and overall utility in most cases, similar to the results observed in the MNIST evaluation.
@@
E.8 Full Results with Standard Deviation
Continuing from Sec. 5.2 and Sec. 5.4, we show full results with standard deviations. Table 11 and Table 12 show the full results of Table 4, demonstrating the stable fairness improvements of PFGuard while preserving utility.